OpenStudy (anonymous):

How do you solve the following two equations by factoring? 1) x² - 2x - 15 = 0 2) 2x² - 3x + 1 = 0 Thank you :)

OpenStudy (anonymous):

You can use the quadratic formula when you get you X's turn them into factors. Do you understand or do you need me to show you?

OpenStudy (anonymous):

Well I'm struggling so I'd like you to show me please. :) @kelsey1

OpenStudy (anonymous):

\[x^2 - 2x - 15 = 0\] This factors into \[(x-5)(x+3) = 0\]. So \[x=5\] or \[x=-3\]

OpenStudy (anonymous):

okay no problem I'll show you the steps!

OpenStudy (anonymous):

Awesome thank you very much! I was stuck on the factoring part. @rsnoob Any ideas for the second equation? @kelsey1

OpenStudy (anonymous):

1. \[2x2 -3x + 1 = 0\] 2. \[2x2 - 2x - x + 1 = 0\] 3. \[2x(x-1)-1(x-1)=0\] 4.\[( 2x-1) (x-1 ) = 0\] 5. \[X = 1/2\] or \[x=1\]

OpenStudy (anonymous):